OpenStudy (anonymous):

find the exact value of cos(pi/12)

OpenStudy (anonymous):

Use a calculator.

OpenStudy (anonymous):

can't use a calculator, pi is a non repeating number, so a decimal wouldn't give and Exact value.

OpenStudy (anonymous):

i have to show the work

OpenStudy (anonymous):

hint, remember the unit circle.

OpenStudy (campbell_st):

use the difference of 2 angles \[\cos(A - B)\] where A = pi/3 and B = pi/4 both the values A and B have exact values. So after you have written the difference in expanded form, make the approriate substitutions.

OpenStudy (anonymous):

the answer is \[\frac{ \sqrt{6} }{ 4 } + \frac{ \sqrt{2} }{ 4 }\] but i need to know how to get there \[\frac{ \sqrt{6} }{ 4 } + \frac{ \sqrt{2} }{ 4 }\]

OpenStudy (campbell_st):

hope it makes sense.

OpenStudy (anonymous):

not really :/

OpenStudy (campbell_st):

ok... do you know the expansion for cos(A - B)..?

OpenStudy (anonymous):

no... im taking precal but i really don't get anything

OpenStudy (campbell_st):

the reason you need it is because \[\frac{\pi}{3} - \frac{\pi}{4} = \frac{\pi}{12}\]

OpenStudy (campbell_st):

well this is the expansion \[\cos(A -B) = \cos(A)\cos(B) - \sin(A)\sin(B)\] so if \[A = \frac{\pi}{3}...and.... B=\frac{\pi}{4}\] you'll have \[\cos(\frac{\pi}{12})=\cos(\frac{\pi}{3} - \frac{\pi}{4}) = \cos(\frac{\pi}{3})\cos(\frac{\pi}{4}) - \sin(\frac{\pi}{3})\sin(\frac{\pi}{4})\] now you have trig ratios and angles that have exact values... make the necessary substitution... and then evaluate for the answer

OpenStudy (campbell_st):

opps should be a + between the terms not a negative \[\cos(\frac{\pi}{3})\cos(\frac{\pi}{4} )+ \sin(\frac{\pi}{3})\sin(\frac{\pi}{4})\]

OpenStudy (anonymous):

[0.35+\sqrt{6}/4\]